A ferris wheel is 30 meters in diameter and boarded from a platform that is 2 meters above the ground. The six o'clock position on the ferris wheel is level with the loading platform. The wheel completes 1 full revolution in 8 minutes. The function h = f(t) gives your height in meters above the ground t minutes after the wheel begins to turn.

What is the Amplitude?  meters
What is the Midline? y =  meters
What is the Period? y =  minutes
How High are you off of the ground after 4 minutes?  meters
Answer #1

Answer :

Given data is :

Diameter = 30 meters

Time period = 8 minutes

No.of meters above the ground = 2

a)Formula to calculate Amplitude is :

A = D / 2

Where D = Diameter

Therefore,

A = 30 / 2 = 15

A = 15 meters

b)Formula to calculate Mid-line is : y

Mid line = A + DG

Where A = Amplitude

G = Distance from ground

Mid line y = 15 + 2 = 17

y = 17 meters

c)Period y = Time taken to complete the 1 revolution = 8 minutes

d)Formula to calculate after 4 minutes is : A + Mid line

After 4 minutes = 15 + 17 = 32

High of the ground after 4 minutes = 32

Given function

Generally

Where

C = 17